Mir Status Report 10 April 1997

Date: 10 April 1997 / Mission Day 89
Mir Station System/Subsystem Activities: IFM's (entire crew)

Mir Systems

1. Jerry and Vasily (Tsibliev) completed removal of the leaking heat exchanger in Kvant, while Sasha (Lazutkin) worked on the Elektron filter bypass in Kvant II.

2. The pressure check of the Kvant VGK cooling loop showed the pressure decreased from 900 mmHg to 770 mmHg in 30 minutes. By the next pass, the crew had repressurized the loop to 898 mmHg, and by 20:01, had held pressure for over one hour. All heat exchangers are still clean and dry.

3. The Elektron in Kvant II was switched on, but then automatically shut down after 9 minutes. A malfunctioning BVKD pressure gauge was also detected. The crew was told to repeat the IFM procedure. It was suggested a new BVKD pressure gauge be added to the STS-84 manifest. At 20:02 DMT, the crew indicated they had activated the Elektron three additional times per the ground instructions and the unit worked for one minute, two minutes, and twelve minutes, respectively, before going into 'auto shutdown'. The ground said telemetry indicated a bad control unit and instructed the crew to swap the unit with the unit from the Kvant Elektron.

4. The Condensate Removal System (CRS) cooling fan in the Soyuz was showing 'negative function.' Currently, all Mir humidity removal is being performed by the Soyuz CRS. The ground is currently writing an IFM r/g for this.

5. According to the most recent 'Mir Onboard Systems and Consumables Status' received today, there are 14 unused LiOH canisters onboard as of yesterday.

Environmental Parameters (Information from 9 April)

Parameter                                  Min.   Max.   Nominal Range 

Pressure (mmHg) Base Block 745 758 660 - 850
ppO2 (mmHg) 160 164 140 - 200
ppCO2 (mmHg) 5.8 6.4 up to 8.0
ppH20 (mmHg) 15.7 16.2 8 - 14
Humidity (percentage) 53 55 30 - 75
Temperature (0C) base block (BB) 28.5 28.9 18 - 28
Temp (0C) docking/transfer module 21.9 22.9 18 - 28
Temp (0C) work area Krystall 25.2 25.5 18 - 28
Temp (0C) sleep area Krystall 26.0 26.2 18 - 28
Temp (0C) Spektr 26.3 26.5 18 - 28
Temp (0C) Kvant II 20.4 21.2 18 - 28


Planning

MD 90: IFM's (entire crew) MD 91: Condensate Presence Check of Mir (Mir 23); Wet Cleanup (entire crew) A/G Audio with Friends in celebration of Cosmonauts Day (entire crew); Rest (entire crew)

MD 92: KOB-2 Loop Leak Repair (Mir 23)

  • submit to reddit


Loading



MILITARY SPACE USA, 11TH JUNE TO 12TH JUNE 2019, LOS ANGELES, USA
49th International Conference on Environmental Systems - ICES 2019




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on April 10, 1997 12:30 PM.

Shuttle Mir Science Program Status Report 3-9 April 1997 was the previous entry in this blog.

Mir Status Report 11 April 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.